OpenStudy (anonymous):

Find the sum. 1 + 3 + 5 + ... + 1743 Express the sum using summation notation. 3 + 12 + 27 + . . . + 75

OpenStudy (anonymous):

\[Sn=(a1+an).n/2\] an=a1+(n-1).r

OpenStudy (anonymous):

Can you find n from the formula?

OpenStudy (anonymous):

yes put an as 1743

OpenStudy (anonymous):

r is 2

OpenStudy (campbell_st):

this is an arithmetic series where you know the 1st and last terms and the common difference so find the number of terms by using \[T _{n} = a + ( n - 1) d\] 1743 = 1 + ( n - 1) 2 1743 = 2n - 1 1744 = 2n n = 872 to find the sum use \[s _{n}= n/2[ a + l]\] a = 1st term l = last term \[s _{872}=872/2[1 + 1744]\] \[s _{872}= 760384\]

OpenStudy (anonymous):

My S_872 = 764, 744

OpenStudy (anonymous):

Oh, n = 877

OpenStudy (campbell_st):

oops \[s _{872}= 872/2[1+1743] = 760384\]

OpenStudy (anonymous):

Campell's right! My manual calculation is off :(

OpenStudy (campbell_st):

lol... the problem we all have